The payroll of YellowCard Company for September 2013 is as follows. Total payroll was $470,400, of which $153,700 is exempt from Social Security tax because it represented amounts paid in excess of $106,800 to certain employees. The amount paid to employees in excess of $7,000 was $391,700. Income taxes in the amount of $88,780 were withheld, as was $8,830 in union dues. The state unemployment tax is 3.5%, but YellowCard Company is allowed a credit of 2.3% by the state for its unemployment experience. Also, assume that the current FICA tax is 7.65% on an employee’s wages to $106,800 and 1.45% in excess of $106,800. No employee for YellowCard makes more than $125,000. The federal unemployment tax rate is 0.8% after state credit.

Prepare the necessary journal entries if (a) the wages and salaries paid (5 lines) and (b) the employer payroll taxes are recorded separately (4 lines).

(Round answers to 0 decimal places, e.g. 5,275. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts .Credit account titles are automatically indented when amount is entered. Do not indent manually.) No. Account Titles and Explanation Debit Credit (a) (b)
Answer #1
Debit Credit
Salaries and Wage Expense 470400
       Withholding Taxes payable 88780
       FICA Taxes Payable 26456 =(470400-153700)*7.65%+(153700*1.45%)
       Union Dues Payable 8830
       Cash 346334
Payroll Tax Expense 28030
       FICA Taxes Payable 26456
       FUTA Taxes Payable 630 =(470400-391700)*0.8%
       SUTA Taxes Payable 944 =(470400-391700)*(3.5%-2.3%)